OpenStudy (anonymous):

Suppose you spend $50.00 on party snacks. The guacamole (variable g) costs $8.00 per pound and the chips (variable c) cost $4.50 per bag. Write an equation in standard form that represents the different amounts (variable A) in pounds of guacamole, g, and chips, c, you could buy. If you buy 4 pounds of guacamole, how many bags of chips can you buy?

OpenStudy (anonymous):

if you buy 4 pounds of G and you spend $32 on that take whats left of the Money as in Subtract 32 from 50 and divide that buy $4.50?:) for Chips? and let me know how many bags you can buy:)

OpenStudy (wolf1728):

(50 - $8 * pounds of guacomole) / 4.50 = bags of potato chips If you buy 4 pounds of guacomole ($50 - $8*4) / 4.50 = 4 bags of potato chips

OpenStudy (wolf1728):

($50 -(chip bags * 4.50)) / 8 = pounds of guacomole If you buy 5 bags of chips ($50 -(5 * 4.50)) / 8 = pounds of guacomole You could buy 27.50 / 8 = pounds of guacomole = 3.44 pounds of guacomole